he mean (μ) of the scale is 55 and the standard deviation (σ) is 14. Assuming that the scores are normally distributed, what is the PROBABILITY that a score falls below 39?
Answer by stanbon(75887)

he mean (μ) of the scale is 55 and the standard deviation (σ) is 14. Assuming that the scores are normally distributed, what is the PROBABILITY that a score falls below 39?
----------------
z(39) = (39-55)/14 = -1.1429
===================================
P(x < 39) = P(z < -1.1429) = 0.1265
==========
Cheers,
Stan H.